OpenStudy (anonymous):

. In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40.

OpenStudy (lucaz):

x-int is the x value when y=0, subtitute y by 0 and solve for x

OpenStudy (lucaz):

it's a quadratic, so 3 times -40 = -120, this is the product of two numbers that aadd up to 19x -5 times 24 = -120 then the equation is 3x^2+(24x-5x)-40=0 3x(x+8)-5(x+8)=0 (x+8)(3x-5)=0, the x-int are the roots x+8=0 -> x=-8 3x-5=0 -> x=5/3